OpenStudy (anonymous):

MEDAl + fan!!!!! Please help 3x x-2 ---- X ----- x2-4 x3 Thank you! :)

OpenStudy (anonymous):

what do u want done here?

OpenStudy (anonymous):

to simplify :)

OpenStudy (anonymous):

hmmm hints x2-4 = (x-2)(x+2) x3 = x(x2) can you continue?

OpenStudy (anonymous):

um wailt a sec

OpenStudy (anonymous):

k

OpenStudy (anonymous):

ohhh so u simplify the left denominator then cancel with right numerator then what elese can be done?

OpenStudy (anonymous):

3 / (x+2)x2 is what I get

OpenStudy (anonymous):

removed an x from left numer. and right denom.

OpenStudy (anonymous):

kk thanks :)

OpenStudy (anonymous):

yw